in how many ways 7 boys sit in a round table so that 2 particulars boys may sit together?
Answer by stanbon(75887)

in how many ways 7 boys sit in a round table so that 2 particulars boys may sit together?
----
The 2 boys can be arranged in 2 ways; they form a unit.
So now you have 6 units.
At a round table those 6 can be arranged is 5! ways.
--------------
Total # of arrangements = 2*5! = 2*120 = 240 ways
=======================
Cheers,
Stan H.